# A nonparametric approach to practical identifiability of nonlinear mixed effects models

### v1

@@ Abstract
Mathematical modelling is a widely used approach to understand and interpret clinical trial data. This modelling typically involves fitting mechanistic mathematical models to data from individual trial participants. Despite the widespread adoption of this individual-based fitting, it is becoming increasingly common to take a hierarchical approach to parameter estimation, where modellers characterize the population parameter distributions, rather than considering each individual independently. This hierarchical parameter estimation is standard in pharmacometric modelling. However, many of the existing techniques for parameter identifiability do not immediately translate from the individual-based fitting to the hierarchical setting. Here, we propose a nonparametric approach to study practical identifiability within a hierarchical parameter estimation framework. We focus on the commonly used nonlinear mixed effects framework and investigate two well-studied examples from the pharmacometrics and viral dynamics literature to illustrate the potential utility of our approach.
@@ 1 Introduction
Many recent advances in parameter identifiability have are focused on individual-level dynamics and do not immediately extend to hierarchical modelling [60, 59]. Consequently, there has been recent interest in extending existing identifiability techniques to a hierarchical modelling framework [67, 33]. For example, Lavielle and Aarons, [40] illustrated how models which are unidentifiable on the individual level may become identifiable at the population level, with examples drawn from pharmacokinetic modelling. In similar work, Duchesne et al., [22] leveraged the parametric representation of population parameter distributions to facilitate model reduction in an in vitro model of erythropoiesis, while El Messaoudi et al., [23] used a simulation-based approach to re-estimate the known parametric population distributions from simulated data. In each of these cases, the identifiability of the NLME model was assessed via estimates of the parametric representation of the population parameter distribution (for example, estimating the mean and variance of a lognormal distribution). This parametric perspective on identifiability of NLME models is a natural extension of the familiar definition of identifiability for non-hierarchical models [33, 40]. However, this approach relies on a priori fixing a parametric representation of the population distribution and implicitly assumes that the individual participants can precisely determine the underlying patient population. Here, we propose a nonparametric approach to perform practical identifiability analysis of NLME models.
@@ 2.1 Friberg model of chemotherapy-induced neutropenia
Neutrophils are the most common white blood cell in humans and an important part of the immune system. Neutropenia, defined as circulating neutrophil cells below a critical level, is a common dose-limiting toxicity of anti-cancer treatments [10, 17, 25]. Friberg et al., [25] developed a semi-mechanistic transit compartmental model of neutrophil production that has become the gold-standard framework for modelling hematopoietic toxicity [17]. This modelling framework has been extended beyond neutrophils to include other hematopoietic lineages and has been used throughout the pharmaceutical sciences [17, 71, 57, 5]. The Friberg et al., [25] model includes the dynamics of proliferating neutrophil precursors ($P(t)$), three compartments representing maturation stages in the bone marrow ($T_{1,2,3}(t)$), and a circulating neutrophil concentration that is typically fit to clinical neutrophil measurements ($N(t)$). The Friberg model is given by
$\frac{\textrm{d}}{\textrm{dt}}P(t)&=\left[(1-E_{drug}(t))\left(\frac{N_{0}}{N(t)}\right)^{\gamma}-1\right]k_{prol}P(t)\\ \frac{\textrm{d}}{\textrm{dt}}T_{1}(t)&=k_{prol}P(t)-k_{tr}T_{1}(t)\\ \frac{\textrm{d}}{\textrm{dt}}T_{i}(t)&=k_{tr}T_{i-1}(t)-k_{tr}T_{i}(t)\quad\textrm{for}\quad i=2,3\\ \frac{\textrm{d}}{\textrm{dt}}N(t)&=k_{tr}T_{3}(t)-k_{circ}N(t).\end{aligned}\right\}$ (1)
In Eq. (1), $k_{prol}$ is the proliferation rate of neutrophil precursors. The proliferation of these cells is modulated by the ratio $N_{0}/N(t)$. This feedback from circulating neutrophils to proliferating neutrophil precursors acts as a cipher for the cytokine effects, and $\gamma$ controls the strength of this feedback. Neutrophil precursors in the bone marrow progress through the maturation compartments with rate $k_{tr}$. Finally, circulating neutrophils are cleared at a per capita rate $k_{circ}.$ As is commonly done [25, 5], we take the initial conditions of Eq. (1) as
$N(0)=N_{0},\quad T_{1,2,3}(0)=\frac{k_{circ}N(0)}{k_{tr}},\quad\textrm{and}\quad P(0)=\frac{k_{circ}N(0)}{k_{prol}}.$
@@
$\frac{\textrm{d}}{\textrm{dt}}C_{p}(t)&=I_{c}(t)+k_{fp}C_{f}(t)+k_{sl1p}C_{sl1}(t)-(k_{pf}+k_{psl1}+k_{cl})C_{p}(t)\\ \frac{\textrm{d}}{\textrm{dt}}C_{f}(t)&=k_{pf}C_{p}(t)k_{sl2f}C_{sl2}(t)-(k_{fp}+k_{fsl2})C_{f}(t)\\ \frac{\textrm{d}}{\textrm{dt}}C_{sl1}(t)&=k_{psl1}C_{p}(t)-k_{sl1p}C_{sl1}(t)\\ \frac{\textrm{d}}{\textrm{dt}}C_{sl2}(t)&=k_{psl2}C_{p}(t)-k_{sl2f}C_{sl2}(t).\end{aligned}\right\}$ (2)
@@ 2.1.2 Generating a virtual population and synthetic data
We generated a virtual population by sampling from the population parameter distributions for docetaxel treatment reported in Table 4 of Friberg et al., [25]. The parameters defining the neutrophil dynamics and antiproliferative effects were taken from Friberg et al., [25] and are therefore self-consistent. We included inter-individual variability on model parameters corresponding to the baseline neutrophil concentration, $N_{0}$, the $EC_{50}$, the maturation rate $k_{tr}$, and, as we are not enforcing $k_{tr}=k_{prol},$ we include inter-individual variability on $k_{prol}$. We assumed that all parameters with inter-individual variability followed a log-normal distribution at the population level. Here, we follow Friberg et al., [25] and log-transform the inter-individual variability sampled from Normal distributions. The mean and variance of these population distributions are given in Table 1.
@@
Table 1: Population parameters used to generate simulated data using the Friberg neutropenia model. A standard deviation of N/A indicates that the given parameter was fixed across the entire virtual population. These parameter values are taken from Table 4 of Friberg et al., [25]. Consistent with the parameter estimation in Friberg et al., [25], we sample from normal distributions before transforming the parameter value to the linear scale.
We utilized the population PK parameter estimates from Pérez-Ruixo et al., [53] to parameterize the PK model in Eq. (2). As we are primarily interested in the identifiability of the Friberg model of chemotherapy-induced neutropenia, we do not include any inter-individual variability on the PK parameters. We generated synthetic circulating neutrophil concentrations by sampling $N=15$ virtual participants from the known population parameter distributions in Table 1. For each virtual participant, we simulate the Friberg model in Eq. (1) for 65 days with Zalypsis administrations on days 0, 21, 42, and 63. We measure the corresponding neutrophil concentration on days $t_{j}=0,3,6,9,...,63$.
As mentioned above, the Friberg model is structurally unidentifiable at the individual level. We wish to test whether simultaneously fitting multiple individuals receiving the same treatment will allow for accurate estimation of the population parameter distributions, even if the underlying individual model is unidentifiable. We therefore consider the idealized situation, where the model is perfectly specified to the experimental data.
@@ 2.1.3 Parameterization of the Friberg model
To fit the synthetic neutrophil data, we first fixed the parameters defining the PK model in Eq. (2) to existing population parameter estimates [19, 10]. Moreover, we fixed the maximal anti-proliferative effect of chemotherapy as $E_{max}=1$. We further assumed that the clearance rate of circulating neutrophils, $k_{circ},$ and the strength of feedback from circulation to proliferation, $\gamma$, are approximately equal between participants in this simulated clinical trial. Thus, we only estimated $k_{circ}$ and $\gamma$ at the population level without inter-individual variability.
We allowed the remaining parameters, $k_{prol},N_{0},EC_{50},$ and $k_{tr}$ to have inter-individual variability. We fit the Friberg model (1) to the synthetic data using the nonlinear mixed-effects framework implemented in Monolix using the stochastic approximation expectation maximization algorithm (SAEM) [39, 63, 38, 20]. We assumed that the model parameters with inter-individual variability, $\phi=[k_{prol},N_{0},EC_{50},k_{tr}]$, follow a log-normal distribution at the population level. The population parameter distribution is defined by
@@
where $\varphi_{i}$ is the mean population estimate and $\psi_{i}\sim\mathcal{N}(0,\omega_{i}^{2})$ captures the inter-individual variability for each parameter [63, 39].
We fit the Friberg model in Eq. (1) to the synthetic neutrophil data sampled at 222222 distinct time points $t_{j}=0,3,6,9,...,63$ for the $N=15$ virtual participants. We denote the synthetic data from the $i$th virtual participant at time $t_{j}$ by $y_{i,j}$. Now, let $\theta_{i}$ be the model parameters for the $i$th virtual participant and $f(t_{j},\theta_{i})$ be the simulated neutrophil concentration, $N(t_{j}),$ for model parameters $\theta_{i}$. For each participant, the statistical model is therefore given by
$y_{i,j}=f(t_{i,j},\theta_{i})+g(f(t_{i,j},\theta_{i}),\omega)e_{i,j}.$
In this framework, the error model is given by $g(f(t_{i,j},\theta_{i}),\omega)e_{i,j}$, where $e_{i,j}$ is the residual error of participant $i$ at measurement $j$, and $\omega$ is a vector of error parameters. The residual error, $e_{i,j}$, is assumed to follow a normal distribution with mean 0 and variance 111, so the error function $g(f(t_{ijk},\theta_{i}),\omega)$ determines the standard deviation of the residual error model [39]. For the synthetic neutrophil concentrations, we consider a proportional error model, so
$g(f(t_{i,j},\theta_{i}),\omega)=b[f(t_{i,j},\theta_{i})],$
@@ 2.2 Standard viral dynamics model
$\frac{\textrm{d}}{\textrm{dt}}T(t)&=\lambda-\beta V(t)T(t)-d_{T}T(t),\\ \frac{\textrm{d}}{\textrm{dt}}I(t)&=\beta V(t)T(t)-\delta I(t),\\ \frac{\textrm{d}}{\textrm{dt}}V(t)&=pI(t)-cV(t).\end{aligned}\right\}$ (3)
@@ 2.2.2 Generating a virtual population and synthetic data
Here, we consider the dynamics of the concentrations of circulating viral loads after acute HIV-1 infection. To generate synthetic viral load data, we sampled model parameters from population distributions to generate virtual patients. We did not fit these population-level parameter values to clinical data. However, they are consistent with estimates in the literature [15, 54, 31] and are given in Table 2.
@@
To determine the initial conditions of the model, we recall that we are considering acute infection. Consequently, we assume that the baseline concentration of uninfected CD4 T-cells is roughly unchanged from HIV-negative individuals. We assume that $\log(T(0))$ follows a normal distribution with mean $\mu_{T(0)}=1.5\times 10^{6}$cells/mL and standard deviation $\sigma_{T(0)}=0.45$. We consider viral load data beginning with the first quantifiable circulating viral load measurement as $V(0).$ As is typical in clinical studies, we consider the $\log_{10}$ transformed circulating viral load. Consequently, we assume that the first quantifiable viral load measurement is normally distributed with mean $1\log_{10}$ copies/mL. This baseline concentration is consistent with the lower limit of quantification of highly sensitive assays [49].
@@
We generate synthetic viral load data by sampling $N=15$ virtual participants from the known population parameter distributions. For each virtual participant, we simulate the viral dynamics model in Eq. (3) for 65 days. We measure the $\log_{10}$ viral load concentrations on days $t_{j}=0,8,12,16,...,64$ and add IID Gaussian measurement error with mean 0 and standard deviation $\sigma=0.1\log_{10}$ copies/mL. Here, we do not consider the idealized scenario, where the synthetic data is exactly generated by the model, of the previous example.
@@ 2.2.3 Parameterization of the standard viral dynamics model
As mentioned, the death rate of uninfected CD4 T-cells is typically set to $d=0.01$/day and the per capita clearance rate of free virus is taken as $c=23$/day. Consequently, we fix these parameters when fitting the synthetic viral load data. We assumed that $V(0)$ represents the first quantifiable viral load measurement. However, this viral load measurement results from an infection event at time $t_{0}<0$. We therefore estimated the concentration of CD4 T-cells, $T_{0}=T(0)$, and viral load, $V_{0}=V(0)$, at time $t=0$.
@@
We fit the standard viral dynamics model Eq. (3) to synthetic HIV viral load that was sampled at 161616 distinct time points, $t_{j}=0,8,12,16,...,64$. For the $i$th virtual participant, the $\log_{10}$ synthetic viral load concentration at time $t_{j}$ is denoted by $y_{i,j}$, while the vector of model parameters for the $i$th individual is given by $\theta_{i}$. As in the case of the Friberg model, the statistical model is given by
$y_{i,j}=\log_{10}(f(t_{i,j},\theta_{i}))+g(f(t_{i,j},\theta_{i}),\omega)e_{i,j},$
where $f(t_{j},\theta{i})$ is the solution $V(t_{j})$ of Eq. (3). There are many common choices for the error function $g(f(t_{ijk},\theta),\omega)$ in viral dynamics modelling [28, 77, 6, 13, 45, 32]. We are estimating the $\log_{10}$ viral load concentration, so we set
$Y_{i,j}=\log_{10}(f(t_{i,j},\phi_{i}))+ae_{i,j}.$

### v2

@@ Abstract
Mathematical modelling is a widely used approach to understand and interpret clinical trial data. This modelling typically involves fitting mechanistic mathematical models to data from individual trial participants. Despite the widespread adoption of this individual-based fitting, it is becoming increasingly common to take a hierarchical approach to parameter estimation, where modellers characterize the population parameter distributions, rather than considering each individual independently. This hierarchical parameter estimation is standard in pharmacometric modelling. However, many of the existing techniques for parameter identifiability do not immediately translate from the individual-based fitting to the hierarchical setting. In this work, we propose a nonparametric approach to study practical identifiability within a hierarchical parameter estimation framework. We focus on the commonly used nonlinear mixed effects framework and investigate two well-studied examples from the pharmacometrics and viral dynamics literature to illustrate the potential utility of our approach.
@@ 1 Introduction
Many recent advances in parameter identifiability have focused on individual-level dynamics and do not immediately extend to hierarchical modelling [70, 69]. Consequently, there has been recent interest in extending existing identifiability techniques to a hierarchical modelling framework [76, 38, 37]. For example, Lavielle and Aarons, [45] illustrated how models which are unidentifiable on the individual level may become identifiable at the population level, with examples drawn from pharmacokinetic modelling. In similar work, Duchesne et al., [23] leveraged the parametric representation of population parameter distributions to facilitate model reduction in an in vitro model of erythropoiesis, while El Messaoudi et al., [24] used a simulation-based approach to re-estimate the known parametric population distributions from simulated data. In each of these cases, the identifiability of the NLME model was assessed via estimates of the parametric representation of the population parameter distribution (for example, estimating the mean and variance of a lognormal distribution). This parametric perspective on identifiability of NLME models is a natural extension of the familiar definition of identifiability for non-hierarchical models [38, 45]. However, this approach relies on a priori fixing a parametric representation of the population distribution and implicitly assumes that the individual participants can precisely determine the underlying patient population. Here, we propose a nonparametric approach to perform practical identifiability analysis of NLME models.
@@ 2.1 Friberg model of chemotherapy-induced neutropenia
Neutrophils are the most common white blood cell in humans and an important part of the immune system. Neutropenia, defined as circulating neutrophil cells below a critical level, is a common dose-limiting toxicity of anti-cancer treatments [10, 18, 26]. Friberg et al., [26] developed a semi-mechanistic transit compartmental model of neutrophil production that has become the gold-standard framework for modelling hematopoietic toxicity [18]. This modelling framework has been extended beyond neutrophils to include other hematopoietic lineages and has been used throughout the pharmaceutical sciences [18, 82, 67, 5]. The Friberg et al., [26] model includes the dynamics of proliferating neutrophil precursors ($P(t)$), three compartments representing maturation stages in the bone marrow ($T_{1,2,3}(t)$), and a circulating neutrophil concentration that is typically fit to clinical neutrophil measurements ($N(t)$). Câmara De Souza et al., [5] showed that the Friberg et al., [26] formulation of this model imposes an implicit relationship between the mean maturation time and the proliferation rate ($k_{prol}$) of neutrophil precursors. However, these two processes are mechanistically distinct. Consequently, Câmara De Souza et al., [5] proposed the following generalization of the Friberg model that does not impose this relationship and is given by
$\left.\begin{aligned} \frac{\textrm{d}}{\textrm{dt}}P(t)&=\left[(1-E_{drug}(t))\left(\frac{N_{0}}{N(t)}\right)^{\gamma}-1\right]k_{prol}P(t)\\ \frac{\textrm{d}}{\textrm{dt}}T_{1}(t)&=k_{prol}P(t)-k_{tr}T_{1}(t)\\ \frac{\textrm{d}}{\textrm{dt}}T_{i}(t)&=k_{tr}T_{i-1}(t)-k_{tr}T_{i}(t)\qquad\textrm{for}\qquad i=2,3\\ \frac{\textrm{d}}{\textrm{dt}}N(t)&=k_{tr}T_{3}(t)-k_{circ}N(t).\end{aligned}\right\}$ (1)
Eq. (1) decouples the proliferation rate $k_{prol}$ from the maturation rate $k_{tr}$, while Friberg et al., [26] makes the modelling assumption that these rates are equal. The proliferation of these cells is modulated by the ratio $N_{0}/N(t)$. This feedback from circulating neutrophils to proliferating neutrophil precursors acts as a cipher for the cytokine effects, and $\gamma$ controls the strength of this feedback. Neutrophil precursors in the bone marrow progress through the maturation compartments with rate $k_{tr}$. Finally, circulating neutrophils are cleared at a per capita rate $k_{circ}.$ As is commonly done [26, 5], we take the initial conditions of Eq. (1) as
$N(0)=N_{0},\qquad T_{1,2,3}(0)=\frac{k_{circ}N(0)}{k_{tr}},\qquad\textrm{and}\qquad P(0)=\frac{k_{circ}N(0)}{k_{prol}}.$
@@
$\left.\begin{aligned} \frac{\textrm{d}}{\textrm{dt}}C_{p}(t)&=I_{c}(t)+k_{fp}C_{f}(t)+k_{sl1p}C_{sl1}(t)-(k_{pf}+k_{psl1}+k_{cl})C_{p}(t)\\ \frac{\textrm{d}}{\textrm{dt}}C_{f}(t)&=k_{pf}C_{p}(t)+k_{sl2f}C_{sl2}(t)-(k_{fp}+k_{fsl2})C_{f}(t)\\ \frac{\textrm{d}}{\textrm{dt}}C_{sl1}(t)&=k_{psl1}C_{p}(t)-k_{sl1p}C_{sl1}(t)\\ \frac{\textrm{d}}{\textrm{dt}}C_{sl2}(t)&=k_{fsl2}C_{f}(t)-k_{sl2f}C_{sl2}(t).\end{aligned}\right\}$ (2)
@@ 2.1.2 Generating a virtual population and synthetic data
We generated a virtual population using the population parameter distributions for docetaxel treatment that are reported in Table 4 of Friberg et al., [26]. The population parameter estimates defining the neutrophil dynamics and antiproliferative effects of the anticancer treatment were taken from Friberg et al., [26] and are therefore self-consistent. We included inter-individual variability on model parameters corresponding to the baseline neutrophil concentration, $N_{0}$, the $EC_{50}$, the maturation rate $k_{tr}$, and, as we are not enforcing $k_{tr}=k_{prol},$ we include inter-individual variability on $k_{prol}$. We assumed that all parameters with inter-individual variability followed a log-normal distribution at the population level. Here, we follow Friberg et al., [26] and log-transform the inter-individual variability sampled from Normal distributions. As our model formulation slightly differs from Friberg et al., [26], we chose the standard deviations of the inter-individual variability to ensure that the simulated neutrophil dynamics reflect the heterogeneity in Figure 1 of Friberg et al., [26]. The mean and variance of these population distributions are given in Table 1.
@@
Table 1: Population parameters used to generate simulated data using the Friberg neutropenia model. A standard deviation of N/A indicates that the given parameter was fixed across the entire virtual population. The population parameter estimates are taken from Table 4 of Friberg et al., [26]. Consistent with the parameter estimation in Friberg et al., [26], we sample from normal distributions before transforming the parameter value to the linear scale.
We utilized the population PK parameter estimates from Pérez-Ruixo et al., [63] to parameterize the PK model in Eq. (2). As we are primarily interested in the identifiability of the Friberg model of chemotherapy-induced neutropenia, we do not include any inter-individual variability on the PK parameters. We generated synthetic circulating neutrophil concentrations for $N=15$ virtual participants, by sampling from the known population parameter distributions in Table 1, reflecting the mean size of phase I trials of docetaxel [15]. For each virtual participant, we simulate the Friberg model in Eq. (1) for 65 days with Zalypsis administrations on days 0, 21, 42, and 63. We measure the corresponding neutrophil concentration on days $t_{j}=0,3,6,9,...,63$.
As mentioned above, the Friberg model is structurally unidentifiable at the individual level. We wish to test whether simultaneously fitting multiple individuals receiving the same treatment will allow for accurate estimation of the population parameter distributions, even if the underlying individual model is unidentifiable. In what follows, we consider the idealized situation, where the neutrophil measurements are noise-free. This implies that the model is perfectly specified to the experimental data. We can therefore study the identifiability of the model parameters without considering any model misspecification or measurement error [4].
@@ 2.1.3 Parameterization of the Friberg model
To fit the synthetic neutrophil data, we first fixed the parameters defining the PK model in Eq. (2) to existing population parameter estimates [20, 10]. Moreover, we fixed the maximal anti-proliferative effect of chemotherapy as $E_{max}=1$. We further assumed that the clearance rate of circulating neutrophils, $k_{circ},$ and the strength of feedback from circulation to proliferation, $\gamma$, are approximately equal between participants in this simulated clinical trial. Thus, we estimated $k_{circ}$ and $\gamma$ at the population level without inter-individual variability. We allowed the remaining parameters, $k_{prol},N_{0},EC_{50},$ and $k_{tr}$ to have inter-individual variability.
We fit the Friberg model (1) to the synthetic data using the nonlinear mixed-effects framework implemented in Monolix using the stochastic approximation expectation maximization algorithm (SAEM) [44, 78, 43, 21]. We assumed that the model parameters with inter-individual variability, $\phi=[k_{prol},N_{0},EC_{50},k_{tr}]$, follow a log-normal distribution at the population level. The population parameter distribution is defined by
@@
where $\varphi_{i}$ is the population estimate, $\omega$ is the standard deviation of the random effects, and $\psi_{i}\sim\mathcal{N}(0,\omega_{i}^{2})$ captures the inter-individual variability for each parameter [78, 44].
We fit the Friberg model in Eq. (1) to the synthetic neutrophil data sampled at $22$ distinct time points $t_{j}=0,3,6,9,...,63$ for the $N=15$ virtual participants. We denote the synthetic data from the $i$th virtual participant at time $t_{j}$ by $y_{i,j}$. Now, let $\theta_{i}$ be the model parameters for the $i$th virtual participant and $f(t_{j},\theta_{i})$ be the simulated neutrophil concentration, $N(t_{j}),$ for model parameters $\theta_{i}$. For each participant, the statistical model is therefore given by
$y_{i,j}=f(t_{j},\theta_{i})+g(f(t_{j},\theta_{i}),\omega)e_{i,j}.$
In this framework, the error model is given by $g(f(t_{j},\theta_{i}),\omega)e_{i,j}$, where $e_{i,j}$ is the residual error of participant $i$ at measurement $j$, and $\omega$ is a vector of error parameters. The residual error, $e_{i,j}$, is assumed to follow a normal distribution with mean 0 and variance $1$, so the error function $g(f(t_{j},\theta_{i}),\omega)$ determines the standard deviation of the residual error model [44]. For the synthetic neutrophil concentrations, we consider a proportional error model, so
$g(f(t_{j},\theta_{i}),\omega)=b[f(t_{j},\theta_{i})],$
@@ 2.2 Standard viral dynamics model
$\left.\begin{aligned} \frac{\textrm{d}}{\textrm{dt}}T(t)&=\lambda-\beta V(t)T(t)-d_{T}T(t),\\ \frac{\textrm{d}}{\textrm{dt}}I(t)&=\beta V(t)T(t)-\delta I(t),\\ \frac{\textrm{d}}{\textrm{dt}}V(t)&=pI(t)-cV(t).\end{aligned}\right\}$ (3)
@@ 2.2.2 Generating a virtual population and synthetic data
Here, we consider the dynamics of the concentrations of circulating viral loads after acute HIV-1 infection. To generate synthetic viral load data, we sampled model parameters from population distributions to generate virtual patients. We did not obtain these population-level parameters by fitting the model to clinical data. However, these population values are broadly consistent with estimates in the literature [16, 64, 34, 49, 12] and are given in Table 2.
@@
To determine the initial conditions of the model, we recall that we are considering acute infection. We consider viral load data beginning with the first quantifiable circulating viral load measurement as $V(0).$ As is typical when modelling viral loads measured in clinical studies, we consider the $\log_{10}$ transformed circulating viral load [11, 35, 71, 64, 84, 72]. Consequently, we assume that the first quantifiable viral load measurement is normally distributed with mean $1\log_{10}$ copies/mL. This baseline concentration is consistent with the lower limit of quantification of highly sensitive assays [59].
@@
Once again, we considered a virtual population reflecting the typical size of phase I trials of anti-HIV antivirals [84, 54, 81]. We generate synthetic viral load data by sampling $N=15$ virtual participants from the known population parameter distributions. For each virtual participant, we simulate the viral dynamics model in Eq. (3) for 65 days. We measure the $\log_{10}$ viral load concentrations on days $t_{j}=0,8,12,16,...,64$ and add IID Gaussian measurement error with mean 0 and standard deviation $\sigma=0.1\log_{10}$ copies/mL. Here, unlike the previous example, we consider the more realistic scenario, where the viral load measurements are not noise-free. Consequently, we study the identifiability of the standard viral dynamics model in the case where there is potential model misspecification or measurement error.
@@ 2.2.3 Parameterization of the standard viral dynamics model
As mentioned, the death rate of uninfected CD4 T-cells is typically set to d=0.01d=0.01/day and the per capita clearance rate of free virus is taken as c=23c=23/day. Consequently, we fix these parameters when fitting the synthetic viral load data. We assumed that $V(0)$ represents the first quantifiable viral load measurement. However, this viral load measurement results from an infection event at time $t_{0}<0$. We therefore estimated the concentration of CD4 T-cells, $T_{0}=T(0)$, and viral load, $V_{0}=V(0)$, at time $t=0$.
@@
We fit the standard viral dynamics model Eq. (3) to synthetic HIV viral load that was sampled at $16$ distinct time points, $t_{j}=0,8,12,16,...,64$. For the $i$th virtual participant, the $\log_{10}$ synthetic viral load concentration at time $t_{j}$ is denoted by $y_{i,j}$, while the vector of model parameters for the $i$th individual is given by $\theta_{i}$. As in the case of the Friberg model, the statistical model is given by
$y_{i,j}=\log_{10}(f(t_{j},\theta_{i}))+g(f(t_{j},\theta_{i}),\omega)e_{i,j},$
where $f(t_{j},\theta{i})$ is the solution $V(t_{j})$ of Eq. (3). There are many common choices for the error function $g(f(t_{j},\theta),\omega)$ in viral dynamics modelling [30, 89, 6, 14, 53, 36]. We are estimating the $\log_{10}$ viral load concentration, so we set
$y_{i,j}=\log_{10}(f(t_{j},\phi_{i}))+ae_{i,j}.$